• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Strauch

    Log entry time 07:59:10 on September 23,1999

    Entry number 27038

    keyword=Owl shift summary

    00:25 : stop run 1763.

    Attention: up to now in the HAC table the alignment values (Flr Mrk and Vernier) were not set in agreement with the actual spectrometer settings. Old values for the electron Flr Mrk = 36.0, Vernier = -41.0 and for the hadron Flr Mrk = 54.5, Vernier = -17. New and actual settings: Electron Flr Mrk = 45.5, Vernier = -7 and for the hadron Flr Mrk = 46.5, Vernier = 2.

    00:25 : change to 15 cm dummy target.

    MCC has problems to deliver more than 6 uA. They try to increase it (was promised to take 10 min).

    00:51 : dummy target runs. Runs 1764 and 1765 with bad beam position and very
    unstable beam.

    00:59 : MCC goes back to 6.4 uA unable to keep the current at 9 uA.
    01:00 : 10 min dummy target run 1766.
    01:10 : change to 12C target
    01:16 : 5 min 12C target run 1767

    01:21 : start configuration change: p_e = 0.763, th_e = 56.536, p_h = 1.0063, th_h = 39.215, fpp: 3" + 6"
    01:51 : 5 min 12C target run 1768
    01:57 : change to 15 cm dummy target
    02:00 : unstable beam / no beam
    02:09 : 12 min dummy target run 1769
    02:21 : change to LH2 target
    02:26 : 27 min H target run 1770, several trips during the run
    02:53 : change to radiator in (6 %)
    03:01 : start radiator in run 1771, bad beam pos. (offset > 1mm) at the end of the run.
    03:11 : start LH2, radiator in run 1772
    03:34 : start LH2, radiator in run 1773
    03:52 : start LH2, radiator in run 1774
    04:10 : change to 15 cm dummy target
    04:16 : start dummy target run 1775 with 4.7 uA, the max. of what MCC can deliver

    04:35 : start configuration change: p_e = 0.661, th_e = 69.54, p_h = 1.1676, th_h
    = 32.044, fpp: 1.5" + 3" + 6"
    05:03 : start 15 cm dummy target run 1776
    05:15 : change to radiator out
    05:18 : start 15 cm dummy target run 1777
    05:35 : change to 12C target
    05:39 : start 12C target run 1778
    05:47 : change to LH2 target
    05:55 : start LH2 target run 1779
    06:08 : MCC tries to increase the beam current
    06:08 : start LH2 target run 1780, beam current changing
    06:21 : accel. tuning without data taking
    06:30 : start LH2 target run 1781
    06:44 : change to radiator in (6 %)
    06:48 : start LH2 target run 1782
    07:26 : start LH2 target run 1783